1 akakakakkakakakakakkakakaka Solve the problem 4log259

Grade:9

1 Answers

Arun
5 years ago
Dear student
 
It can be written as
 
(4*9) log 25
 
36 log (10/2) = 36 log 10 - 36 log 2 = 36 (1-log 2)
 
Hope it helps
 
 
Regards
Arun